Operation Summary for October 10, 2011 - October 12, 2011 |
Monday October 10 The day shift began with Operations monitoring protons in the Recycler (8.7E10), with the Linac, Booster, and Main Injector (MI) operating normally, MiniBooNE experiment with off due to Horn problems, and with the NuMI and experiment taking beam. During the day shift, at 8:31 AM, Operations turned off Linac beam so that Linac personnel could work on a RF station (LRF5); operators loaded an energy conservation timeline into the Time Line Generator. At 9:12 AM, the MiniBooNE Horn tripped off on a ground fault indication. Linac personnel completed their work on LRF5 and resumed sending beam at 9:14 AM; MiniBooNE resume taking beam at this time. An Antiproton Source (Pbar) person began taking one shots of reverse protons at 10:23 AM. At 10:52 AM, Operations tried to send beam to the Meson FTBF experiment T-992, but a power supply problem (F:MT5E) was holding off beam; a Power Supply person investigated. The Pbar reverse proton studies ended at 11:15 AM; operators began monitoring protons in the Debuncher and Accumulator. The Power Supply person reported at 12:54 PM that he had fixed F:MT5E’s cooling fan. Operations turned off beam to NuMI at 2:54 PM for work on their ventilation system. Operations established beam to T-992 at 3:36 PM. During the evening shift, Operations established beam to NuMI at 4:16 PM. At 7:38 PM, a NuMI BPM VME stopped responding; operators rebooted the VME. Tuesday October 11 The midnight shift began with Operations monitoring protons in the Recycler (6.5E10), and with the NuMI, MiniBooNE, and T-992 experiments taking beam. During the midnight shift, Operations disabled beam to T-992 at 12:19 AM. Operators reported that a MI Lambertson (I:LAM42) power supply tripped off for the first of four times in about three hours due to a ripple; it reset. During the day shift, Operations disabled beam to MiniBooNE at 8:12 AM per request. Operators reported at 8:30 AM that a scheduled MTA power outage for power testing had started. At 9 AM, Neutron Therapy Facility (NTF) personnel began to prepare for treating one patient in the afternoon. The MTA power testing ended at 10:01 AM. Operations turned off beam to NuMI at 11:08 AM. Operations turned off beam to MI so that EE Support personnel could work on I:LAM42. An RF person worked on the MI RF system. Operations turned off beam to Booster at 11:58 AM and operators prepared the area for access. Operators prepared MiniBooNE for an access to inspect their Horn at 12:28 PM; they found a small LCW leak. FESS personnel reported at 1:18 PM that they were changing out the MTA feeder breaker trip unit. Booster came out of access at 1:35 PM. Operations established beam to the MI dump at 2:15 PM. Operations established beam to T-992 at 2:26 PM. FESS reported at 3:03 PM that the MTA breaker work was finished. During the evening shift, Operations established beam to NuMI at 4:28 PM. Wednesday October 12 The midnight shift began with Operations monitoring protons in the Recycler (11.6E10), and with the NuMI, MiniBooNE, and T-992 experiments taking beam. During the midnight shift, Operations disabled beam to T-992 at 2 AM. At 2:45 AM, I:LAM42 tripped off; it reset. At 6:24 AM, operators reported a small power glitch that tripped off the MI south anode power supply, all the Booster RF modulators, LRF1, MIRF1, a Linac Klystron RF station (KRF7), three Linac NTF quadrupole power supplies (L:QPSC3, 5, and 6), the Booster ORBUMP, and the LRF5 vacuum valves went to an intermediate state. Operators reestablished beam at 6:39 AM. Operations Coordinator (OPSCO) Plans for Today The plans for today are keep everything running normally. The OPSCO will allow Booster and Pbar Personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 10/12/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |